OpenStudy (dtan5457):

Name 2 complex fractions that equal to x-3/x+4

myininaya (myininaya):

I hope you mean compound fraction because I know what that means here is a hint \[\frac{x-3}{x+4} =\frac{\frac{x-3}{1}}{\frac{x+4}{1}}\]

OpenStudy (fibonaccichick666):

oh, I was thinking about multiplying by some complex number to get the same result

myininaya (myininaya):

\[\frac{x-3}{x+4} =\frac{\frac{x-3}{a(x)}}{\frac{x+4}{a(x)}} \]

OpenStudy (campbell_st):

what about \[\frac{x}{x + 4} - \frac{3}{x + 4} \] perhaps its to simple.

OpenStudy (dtan5457):

can you explain the a(x) in your formula

myininaya (myininaya):

It is a function of x

myininaya (myininaya):

\[\frac{x-3}{x+4} =\frac{\frac{x-3}{1}}{\frac{x+4}{1}} \cdot \frac{\frac{1}{a(x)}}{\frac{1}{a(x)}} =\frac{\frac{x-3}{a(x)}}{\frac{x+4}{a(x)}}\]

OpenStudy (dtan5457):

im pretty sure a complex fraction has to consist of a fraction on top of a fraction

myininaya (myininaya):

(x-3)/a is a fraction (x+4)/a is a fration

OpenStudy (fibonaccichick666):

I've never heard of a camplex fraction so I think of imaginary numbers since they are called complx

OpenStudy (fibonaccichick666):

Let me clear this up thanks to purplemath http://www.purplemath.com/modules/compfrac.htm